Cory borrows $182 for school trip; agrees to pay parents back monthly; 1st month $2, 2nd $4, $6, $8, etc. How many months to pay back?
Answer by stanbon(75887)

Cory borrows $182 for school trip; agrees to pay parents back monthly; 1st month $2, 2nd $4, $6, $8, etc. How many months to pay back?
-----
An arithmetic series: 2,4,6,8....24,26
---
Equation:
S(n) = (n/2)(1st + last)
182 = (n/2)(2+(n-1)2)
364 = n(2+2n-2)
364 = 2n^2
n^2 = 182
n = 13.49
Round up to n=14
Then n-1 = 13 weeks.
----
Cheers,
Stan H.